I dont understand the following: solve by graphing
x+y=7
2x-y=5
The instructions also say if you solve the system and the variables both add out, you need to tell whether the system consists of lines that coincide (infinite solutions) or parallel lines(no solution)

Answer by rapaljer(4671)

Just add the two equations together, and the y terms subtract out:
x+y=7
2x-y=5
--------------
3x=12
x=4

Substitute back into the first equation:
x+y=7
4+y=7
y=3

Check by substituting both values in the second equation:
2x-y=5
2*(4)-3=5
It checks.

The solution is the point (4,3).

R^2